OpenStudy (anonymous):

The point P(1,1/2) lies on the curve y=x/(1+x). If Q is the point (x,x/1+x), find the slope of the secant line PQ when x+ 0.5

OpenStudy (anonymous):

when x=0.5

OpenStudy (anonymous):

Since we only need to find the slope of secant line PQ, all we need are point P and Q. P is given (1, 1/2). Q = (0.5, 0.5/1.5) = (1/2, 1/3). slope = rise/run = (1/3 - 1/2) / (1/2 - 1) = 1/3